OpenStudy (anonymous):

We just started Logarithms. 2^(2x+1) - 3*2^(x) +1 = 0 Thanks!

OpenStudy (anonymous):

5log(c^2-7)-8log(c^-1-2)

OpenStudy (anonymous):

uhh, care to show working out?

OpenStudy (anonymous):

should i put the i simplify that ir should i find the aproximate value??

OpenStudy (anonymous):

i want to show but im lazy using the equation..

OpenStudy (chaise):

I can put the equation into my calculator as well, he wants working out.

OpenStudy (anonymous):

uhh, care to show working out?

OpenStudy (anonymous):

Can I just do this (2x+1) log 2 - 3*x log 2 +1 =0

OpenStudy (nikvist):

\[2^{2x+1}-3\cdot 2^x+1=0\] \[2\cdot 2^{2x}-3\cdot 2^x+1=0\] \[(2^x-1)\cdot(2\cdot 2^x-1)=0\] \[2^x=1\quad\vee\quad 2^x=1/2\] \[x=0\quad\vee\quad x=-1\]